OpenStudy (anonymous):

What's 8x-24 over 3x-9

OpenStudy (anonymous):

Factor the numerator and denominator, then simplify.

OpenStudy (anonymous):

\[\frac{ 8x-24 }{ 3x-9 }=\frac{ 8(x-3) }{ 3(x-3) }=\frac{ 8 }{ 3 }\]